# GCP: Guarded Collaborative Perception with Spatial-Temporal Aware Malicious Agent Detection

### v1

@@ IV-E Joint Spatial-Temporal Benjamini-Hochberg Test
To jointly utilize spatial and temporal anomaly results for malicious agent detection, we employ the Benjamini-Hochberg (BH) procedure [32, 16]. The BH procedure effectively controls the False Discovery Rate (FDR) - the expected proportion of false positives among all rejected null hypotheses, which is crucial in CP where false accusations can severely impact system reliability. For each CAV $i$ under inspection, we first compute a weighted combination of spatial and temporal scores:
@@
where $\omega_{s}$ and $\omega_{t}$ are learnable weights. The BH procedure controls FDR through a step-up process that adapts its rejection threshold based on the distribution of observed p-values. Unlike traditional single hypothesis testing, the BH procedure maintains FDR control at level $\alpha_{bh}$ even under arbitrary dependencies between tests, making it particularly suitable for CP where agents’ behaviors may be correlated due to shared environmental factors or coordinated attacks. We formulate the hypothesis test =$\mathcal{H}=\{\mathcal{H}_{0},\mathcal{H}_{1}\}$, where $\mathcal{H}_{0}:\mathcal{L}_{ST}\sim P_{normal}$ represents the null hypothesis that the combined score follows the distribution of normal agents, and $\mathcal{H}_{1}:\mathcal{L}_{ST}\nsim P_{normal}$ represents the alternative hypothesis. Since this distribution is typically intractable, we compute conformal p-values using the calibration set $\mathcal{S}$:
@@
where ||$\cdot$|| ⋅ | represents the cardinality of the set. Following the BH procedure, we sort the p-values in ascending order $\hat{p}_{(1)}\leq\hat{p}_{(2)}\leq...\leq\hat{p}_{(m)}$ for $m$ hypothesis tests. Let $j$ be the largest index satisfying:
@@
Input:
  • $\mathbf{O}_{e,i}^{t_{k}}$, low-confidence detections at current frame
  • $\{\mathbf{Y}_{e,i}^{t_{k-j}}\}_{j=1}^{K}$, cached detections from past $K$ frames
@@
2:     $\mathbf{I}_{c}^{t_{k}},\mathbf{I}_{u}^{t_{k}}\leftarrow\{\},\{\}$triangleright BEV Flow Initialization
@@
7:         for $k\leftarrow K-1$ down to 00 do $\triangleright$ Chain Match
@@
17:         if $matched$ then $\triangleright$ Flow Concatenation
@@ V Experiments
In this section, we first describe the experimental setup, including dataset, attack and defense settings, and implementation details. After that, both the quantitative and qualitative results are introduced.
@@ V-A Experimental Setup
  • V2X-Flow Dataset. To facilite the pre-training of LSTM-AE used in our GCP, we construct a V2X-Flow dataset based on V2X-Sim. We separate the annotations of the V2X-Sim dataset to generate the BEV flow of each bounding box using the chain-based BEV flow matching method mentioned in Section IV-C. The dataset split of V2X-Flow is the same as V2X-Sim.
Attack Settings. To simulate realistic attack scenarios in vehicular networks, we consider that malicious agents would adopt stealthy strategies instead of continuously launching attacks. The temporal patterns of such attacks can be modeled with different distributions. Given the total number of collaborative agents $N_{a}$, the time horizon $T$, the number of malicious agents $m$, and an attack ratio $\lambda$ (the proportion of malicious messages to the total messages from all agents over the time horizon), we evaluate three representative time series attack modes. For all modes, the total malicious messages $\lambda N_{a}T$ are distributed among $m$ malicious agents following a truncated normal distribution $\mathcal{N}(\frac{\lambda N_{a}T}{m},\sigma^{2})$ to simulate varying attack capabilities while maintaining reasonable attack intensities:
@@
Implementation Details. Our GCP is implemented with PyTorch. Each agent’s locally captured LiDAR or camera data is first encoded into a BEV feature map with size $16\times 16$ and 512 channels. The encoder network adopts a ResNet-style architecture with 5 layers of feature encoding, with output feature dimensions progressively increasing from 32 to 512 channels while spatial dimensions decrease from 256 to 16. The decoder consists of three convolutional blocks with channel dimensions progressively decreasing from 512 to 64. The detection head follows a two-branch design (classification and regression) with standard convolutional layers, batch normalization, and ReLU activation. It uses an anchor-based detection decoder [38] with multiple anchor sizes per location. The fusion method is V2VNet [22]. The LSTM-AE model consists of 2 LSTM layers with hidden dimension 32. We train the model using Adam optimizer (lr=0.001) with the default history sequence length 5. All experiments are conducted on a server with 2 Intel(R) Xeon(R) Silver 4410Y CPUs, 4 NVIDIA RTX A5000 GPUs, and 512 GB RAM.
Baselines and Evaluation Metrics. We compare our method with two state-of-the-art CP defense baselines: ROBOSAC [15] and MADE [16]. We also include three reference settings: (1) Upper-bound - CP without malicious agents, representing the optimal performance; (2) Lower-bound - ego vehicle’s local perception only; and (3) No defense - CP with malicious agents but without any defense measures. To validate the effectiveness of our spatial-temporal design, we also evaluate two variants of GCP: spatial-only defense (GCP-S) and temporal-only defense (GCP-T). The performance of all methods is evaluated using both accuracy and efficiency metrics: average precision at 0.5 IoU (AP@0.5) and 0.7 IoU (AP@0.7) for detection accuracy and Frames Per Second (FPS) for computational efficiency.
@@ V-B Quantitative Results
TABLE I: Comparative results under different attack methods and modes. We report the AP@0.5 and AP@0.7 under different attack methods and temporal patterns. Attack settings: the number of malicious agents $m=2$; attack ratio $\lambda=0.25$; input/output perturbation budget $\Delta_{i}=\Delta_{o}=0.5$; the number of iterations = 10. Best results excluding upper-bound are in bold, while significant performance drops are marked with $\downarrow$.
@@
MADE [16] Average 77.07 75.75 76.28 74.87 76.61 75.06 64.00$\downarrow$ 54.48$\downarrow$
@@
No Defense Average 36.65$\downarrow$ 35.92$\downarrow$ 17.70$\downarrow$ 15.14$\downarrow$ 38.27$\downarrow$ 37.28$\downarrow$ 54.83$\downarrow$ 45.11$\downarrow$
@@
Comparative results against different attacks. As shown in Table I, we evaluate our proposed GCP against various attack methods and time series modes on the V2X-Sim dataset. First, all defense methods demonstrate varying degrees of effectiveness compared to the no-defense baseline, which suffers severe performance degradation under all attack scenarios, especially for C&W attacks. While baseline methods show some defense capability, with MADE performing slightly better than ROBOSAC against conventional attacks (PGD, C&W, and BIM), they both exhibit significant vulnerability to our proposed BAC attack. Our proposed GCP consistently performs better across all attack scenarios and time series modes. For conventional attacks, GCP outperforms existing methods with consistent improvements (0.47%-0.79% in AP@0.5) across PGD, C&W, and BIM attacks. The advantage becomes more pronounced under our challenging BAC attack, where GCP significantly surpasses MADE and ROBOSAC by 12.64% and 8.53% in AP@0.5, respectively, while maintaining performance close to the upper bound with only 3.88% degradation. Across different time series attack modes (R-Mode, P-Mode, and S-Mode), GCP demonstrates consistent robustness, maintaining stable performance where baseline methods show varying degrees of degradation. These results validate the effectiveness of our spatial-temporal defense mechanism in handling diverse attack patterns in CP systems.
Comparative results with dynamic attack intensities. Table II shows the performance comparison under different perturbation budgets and attack intensities. All methods show performance degradation with more intense attacks ($m=3$, $\lambda=0.5$), where MADE exhibits the largest drops (13.69%-15.68% in AP@0.5) while GCP maintains relatively stable performance with only 5.46%-5.74% degradation. For perturbation budget settings, we observe that larger output perturbation ($\Delta_{o}=0.5$) enables more effective defense than smaller ones ($\Delta_{o}=0.1$). Increasing input perturbation ($\Delta_{i}=1.0$) significantly impacts baseline methods, with ROBOSAC showing 4.56% drop in AP@0.5 even under moderate attacks. In contrast, GCP maintains consistent performance across all settings (less than 0.5% variation in AP@0.5). Overall, our proposed GCP outperforms baselines by 7.12%-18.29% in AP@0.5, with advantages more pronounced under intense attacks and large perturbation budgets. These results demonstrate our spatial-temporal defense mechanism’s strong adaptability and robustness against varying attack parameters.
TABLE II: Comparative performance under different attack intensities. We report the AP@0.5 and AP@0.7 under different input/output perturbation budget ($\Delta_{i}$/$\Delta_{o}$) and attack intensities. Attack settings: moderate attack ($m=2$, $\lambda=0.25$) and intense attack ($m=3$, $\lambda=0.5$); input/output perturbation budget $\Delta_{i}=\Delta_{o}=0.5$. Best results excluding upper-bound are in bold, while significant performance drops are marked with $\downarrow$.
AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7
Upper-bound 80.52 78.65 80.52 78.65 80.52 78.65 80.52 78.65
GCP (Ours) Moderate 75.80 74.96 75.23 74.02 75.58 74.28 75.51 74.69
Intense 70.22 68.57 70.03 68.36 70.07 68.59 69.77 68.20
MADE [16] Moderate 65.63 56.92 65.80 57.26 65.39 55.56 66.93 56.79
Intense 51.94$\downarrow$ 40.87$\downarrow$ 54.22$\downarrow$ 41.38$\downarrow$ 53.21$\downarrow$ 40.61$\downarrow$ 51.04$\downarrow$ 36.25$\downarrow$
ROBOSAC [15] Moderate 68.68 64.82 66.26 62.38 64.12 58.08 65.64 61.42
Intense 60.09 54.88 61.20 55.07 59.08 51.14 61.02 53.87
No Defense Moderate 55.19$\downarrow$ 43.99$\downarrow$ 65.05 55.44 55.92$\downarrow$ 42.91$\downarrow$ 59.54 47.27$\downarrow$
Intense 35.63$\downarrow$ 25.35$\downarrow$ 47.82$\downarrow$ 32.55$\downarrow$ 33.40$\downarrow$ 17.98$\downarrow$ 41.85$\downarrow$ 23.24$\downarrow$
Figure 5: Comparative results of AP under different cached frame length and consecutive KF interpolation times on V2X-Sim Dataset. Attack settings: $m=2$, $\lambda=0.25$; $\Delta_{i}=\Delta_{o}=0.5$. (a)-(d): AP@0.7 results under different attacks and interpolation budgets; (e)-(h): AP@0.7 results under different attacks and cached frame length.
Ablation studies. To evaluate the effectiveness of our proposed GCP, we compare it with its variants that use only spatial or temporal defense mechanisms. As shown in Table III, both components contribute to the overall defense performance but with different strengths against different attacks. For conventional PGD attacks, the spatial component plays a dominant role, with GCP-S achieving comparable performance (77.21% AP@0.5) to the full model (77.54% AP@0.5), while GCP-T shows limited effectiveness (70.16% AP@0.5). However, for our challenging BAC attack, neither component alone can maintain robust performance, with significant drops in GCP-S (7.41% decrease in AP@0.5) and GCP-T (10.63% decrease in AP@0.5) compared to the full model. This demonstrates the necessity of combining both spatial and temporal consistency checks for effective defense against sophisticated attacks in CP systems. Notably, despite using a lighter architecture, GCP-S maintains better performance than ROBOSAC [15] and MADE [16], validating the effectiveness of our CSCLoss design proposed in Section 4.2.
[…]
Quantitative analysis of attack and defense speeds. As shown in Table IV, we analyze computation speeds across different attack and defense methods. For attacks, the reported FPS values represent single-iteration performance. With 3-5 iterations typically sufficient for effective attacks, all methods can enable real-time attacks in low frame rate scenarios (10-15 FPS). Our BAC attack shows slightly lower speed than PGD (22.4 vs 25.3 FPS) due to mask generation but maintains efficiency through a low-frequency mask update strategy (0.5 FPS). On the defense side, while baseline methods MADE and ROBOSAC show higher speeds against BAC attack (82.8 and 44.6 FPS), this actually indicates that their detection mechanisms are bypassed. In contrast, our GCP maintains robust defense speeds (30.2-47.6 FPS) with only 3.88% performance degradation under BAC attack. These results validate that both our attack and defense methods meet the real-time requirements for practical deployment, demonstrating the potential for seamless integration into existing CP systems without compromising efficiency.
TABLE III: Ablation study results. Comparison between full GCP model and its variants (GCP-S and GCP-T). Attack settings: $m=2$, $\lambda=0.25$, $\Delta_{i}=\Delta_{o}=0.5$.
@@
TABLE IV: Quantitative results of attack and defense speed. We report the average computation speed (FPS) under different attack and defense settings. Attack settings: $m=2$, $\lambda=0.25$; $\Delta_{i}=\Delta_{o}=0.5$.
@@
V-C Qualitative Results

### v2

@@ IV-E Joint Spatial-Temporal Benjamini-Hochberg Test
To jointly utilize spatial and temporal anomaly results for malicious agent detection, we employ the Benjamini-Hochberg (BH) procedure [3, 35]. The BH procedure effectively controls the False Discovery Rate (FDR), the expected proportion of false positives among all rejected null hypotheses, which is crucial in CP where false accusations can severely impact system reliability. For each CAV $i$ under inspection, we first compute a weighted combination of spatial and temporal scores:
@@
where $\omega_{s}$ and $\omega_{t}$ are learnable weights. The BH procedure controls FDR through a step-up process that adapts its rejection threshold based on the distribution of observed p-values. Unlike traditional single hypothesis testing, the BH procedure maintains FDR control at level $\alpha_{bh}$ even under arbitrary dependencies between tests, making it particularly suitable for CP where agents’ behaviors may be correlated due to shared environmental factors or coordinated attacks. We formulate the hypothesis test $\mathcal{H}=\{\mathcal{H}_{0},\mathcal{H}_{1}\}$, where $\mathcal{H}_{0}:\mathcal{L}_{ST}\sim P_{normal}$ represents the null hypothesis that the combined score follows the distribution of normal agents, and $\mathcal{H}_{1}:\mathcal{L}_{ST}\nsim P_{normal}$ represents the alternative hypothesis. Since this distribution is typically intractable, we compute conformal p-values using the calibration set $\mathcal{S}$:
@@
where $|\cdot|$ represents the cardinality of the set. Following the BH procedure, we sort the p-values in ascending order $\hat{p}_{(1)}\leq\hat{p}_{(2)}\leq...\leq\hat{p}_{(m)}$ for $m$ hypothesis tests. Let $j$ be the largest index satisfying:
@@
Input: $\mathbf{O}_{e,i}^{t_{k}}$ (low-confidence detections), $\{\mathbf{Y}_{e,i}^{t_{k-j}}\}_{j=1}^{K}$ (cached detections)
@@
2:  $\mathbf{I}_{c}^{t_{k}},\mathbf{I}_{u}^{t_{k}}\leftarrow\{\},\{\}$ $\triangleright$ BEV Flow Initialization
@@
7:    for $k\leftarrow K-1$ down to 0 do $\triangleright$ Chain Match
@@
17:    if matchedmatched then $\triangleright$ BEV Flow Concatenation
@@ V-A Experimental Setup
V2X-Flow Dataset. To facilite the pre-training of LSTM-AE used in our GCP, we construct a V2X-Flow dataset based on V2X-Sim. We separate the annotations of the V2X-Sim dataset to generate the BEV flow of each bounding box using the chain-based BEV flow matching method mentioned in Section IV-C.
Attack Settings. To simulate realistic stealthy attacks, we model temporal patterns where malicious agents attack intermittently. Given $N_{a}$ total agents, $m$ malicious agents, and time horizon $T$, we define the attack ratio $\lambda$ as the proportion of malicious messages in the total traffic. We evaluate three modes where the total malicious load $\lambda N_{a}T$ is distributed among $m$ attackers via a truncated normal distribution $\mathcal{N}(\frac{\lambda N_{a}T}{m},\sigma^{2})$, reflecting varying attack capabilities:
@@
Implementation Details. Our GCP is implemented with PyTorch. Each agent’s locally captured LiDAR or camera data is first encoded into a BEV feature map with size $16\times 16$ and 512 channels. The encoder network adopts a ResNet-style architecture with 5 layers of feature encoding, with output feature dimensions progressively increasing from 32 to 512 channels while spatial dimensions decrease from 256 to 16. The decoder consists of three convolutional blocks with channel dimensions progressively decreasing from 512 to 64. The detection head follows a two-branch design (classification and regression) with standard convolutional layers, batch normalization, and ReLU activation. It uses an anchor-based detection decoder [23] with multiple anchor sizes per location. The fusion method is V2VNet [27]. The LSTM-AE model consists of 2 LSTM layers with hidden dimension 32. We train the model using Adam optimizer (lr=0.001) with the default history sequence length 5. All experiments are conducted on a server with 2 Intel(R) Xeon(R) Silver 4410Y CPUs, a single NVIDIA RTX A5000 GPU, and 512 GB RAM.
Baselines and Evaluation Metrics. We compare our method with state-of-the-art defenses ROBOSAC [17], MADE [35], and CP-Guard [14]. We also introduce a simple Gated Late Fusion as a baseline, which aggregates proposal boxes from all connected agents and filters them using a consistency voting mechanism. Reference settings include Upper-bound (clean CP), Lower-bound (ego-only), and No Defense. Evaluation metrics include Average Precision (AP@0.5, AP@0.7) for accuracy and Frames Per Second (FPS) for efficiency.
@@ V-B Quantitative Results
Benchmark Comparison. Table I evaluates GCP on V2X-Sim. First, while No Defense suffers severe degradation, all defenses show effectiveness. Notably, CP-Guard and Gated Late Fusion are competitive against conventional attacks. However, despite MADE slightly outperforming ROBOSAC, all baselines are vulnerable to our BAC attack. In contrast, GCP excels. For conventional attacks, it achieves consistent improvements (0.47%-0.79% AP@0.5). The advantage is pronounced under BAC attack, where GCP surpasses CP-Guard, Gated Late Fusion, MADE, and ROBOSAC by 9.42%, 6.26%, 12.64%, and 8.53% respectively, maintaining near-upper-bound performance. Across R/P/S-Modes, GCP demonstrates consistent robustness where baselines degrade, validating our spatial-temporal defense against diverse attacks.
TABLE I: Comparative results under different attack methods and modes. Settings: $m=2,\lambda=0.25,\Delta_{i}=\Delta_{o}=0.5$, iterations=10. Bold/Underlined: Best/Second-best (excl. upper-bound). $\downarrow$: Significant drop.
@@
[…]
R 76.12 74.98 75.45 73.88 76.35 74.88 67.78 63.95
P 76.25 74.55 75.92 74.38 76.45 75.38 66.88 64.55
S 77.25 76.18 76.08 75.08 75.55 73.78 66.98 63.15
Gated Late Fusion Avg. 71.18 68.82 70.52 67.98 71.08 68.75 70.38 68.02
R 70.65 68.12 70.58 67.58 70.42 68.08 69.62 67.42
P 71.02 68.64 70.78 68.48 71.05 68.72 70.05 67.52
S 71.95 69.62 70.18 67.72 71.68 69.42 71.48 69.12
[…]
@@
[…]
@@
TABLE II: Performance comparison (AP@0.5 / AP@0.7) under Independent vs. Colluding attacks ($m=2$).
Attack Method Independent Colluding
AP@0.5 AP@0.7 AP@0.5 AP@0.7
PGD No Defense 36.65 35.92 34.35 33.66
GCP (Ours) 77.54 76.51 74.51 73.53
C&W No Defense 17.70 15.14 16.42 14.05
GCP (Ours) 76.81 75.56 73.89 72.68
BIM No Defense 38.27 37.28 35.88 34.91
GCP (Ours) 77.40 76.46 74.65 73.72
BAC No Defense 54.83 45.11 51.92 42.85
GCP (Ours) 76.64 75.64 74.28 73.35
TABLE III: Impact of the number of malicious agents ($m$) on detection performance (under PGD attack).
# Mal. ($m$) AP@0.5 AP@0.7
No Defense GCP (Ours) No Defense GCP (Ours)
0 (Benign) 80.52 78.42 78.65 77.51
1 50.32 78.32 48.42 77.05
2 (Default) 36.65 77.54 35.92 76.51
3 22.11 73.08 21.56 71.88
TABLE IV: Ablation study. GCP vs. variants (GCP-S/T). Settings: $m=2,\lambda=0.25,\Delta_{i}=\Delta_{o}=0.5$.
@@
Robustness under Different Attack Configurations. We first evaluate stability under varying intensities and attacker counts. Table V shows that under intense attacks ($m=3$, $\lambda=0.5$), GCP maintains stability with only 5.6% degradation, significantly outperforming baselines. It also shows consistent performance (<<0.5% variation) across different perturbation budgets. Furthermore, Table III confirms resilience to increasing attacker counts. In benign scenarios ($m=0$), GCP matches the Upper-bound, ensuring minimal impact. As $m$ increases, GCP maintains robust performance (around 77% for $m\leq 2$) while No Defense degrades catastrophically. This confirms GCP’s reliability in highly compromised networks.
Robustness against More Attack Strategies. We further evaluate GCP against two advanced strategies. ❶ Coordinated Attacks: As shown in Table II, even when agents optimize a joint adversarial objective, GCP remains highly effective (AP@0.5 >> 73%). ❷ LiDAR Spoofing Attacks: We implement a physically realistic Ray-Casting attack [34] that injects synthesized points by calculating intersections between LiDAR rays and ghost vehicle meshes. Although this generates plausible raw data perturbations, the resulting ghost boxes are often sparse and intermittent due to optimization constraints. GCP effectively identifies these anomalies, achieving an 84.5% defense success rate (at $K=5$).
TABLE V: Performance under different attack intensities. Settings: Moderate ($m=2,\lambda=0.25$) vs. Intense ($m=3,\lambda=0.5$). Bold/Underlined: Best/Second-best (excl. upper-bound). $\downarrow$: Significant drop.
AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7
Upper-bound 80.52 78.65 80.52 78.65 80.52 78.65 80.52 78.65
GCP (Ours) Moderate 75.80 74.96 75.23 74.02 75.58 74.28 75.51 74.69
Intense 70.22 68.57 70.03 68.36 70.07 68.59 69.77 68.20
CP-Guard Moderate 72.15 68.25 69.88 65.55 67.55 61.25 69.25 64.55
Intense 63.55 57.88 64.55 58.25 62.15 54.25 64.15 56.95
Gated Late Fusion Moderate 69.62 67.42 69.12 66.58 69.42 66.82 69.35 67.18
Intense 64.48 61.65 64.32 61.48 64.35 61.62 64.08 61.32
MADE Moderate 65.63 56.92 65.80 57.26 65.39 55.56 66.93 56.79
[…]
ROBOSAC Moderate 68.68 64.82 66.26 62.38 64.12 58.08 65.64 61.42
Intense 60.09 54.88 61.20 55.07 59.08 51.14 61.02 53.87
[…]
Figure 6: Comparative results under different cached frame length and consecutive KF interpolation times. Attack settings: $m=2$, $\lambda=0.25$; $\Delta_{i}=\Delta_{o}=0.5$. (a)-(d): AP@0.7 under different attacks and interpolation budgets; (e)-(h): AP@0.7 under different attacks and cached frame length.
TABLE VI: Resource consumption and runtime breakdown.
Resource Consumption
Metric No Defense GCP (Ours) Overhead
GPU Memory (MB) 28,020 28,138 +118
GPU Utilization (%) 0.6 0.9 +0.3
CPU Utilization (%) 10.4 6.7 -3.7
Runtime Breakdown
Module Time (ms) Equiv. FPS
Spatial Defense (CSCLoss) 3.2 -
Temporal Defense (Flow Gen.) 1.8 -
Temporal Defense (LSTM-AE) 21.0 -
Total Defense Latency 26.0 38.6
TABLE VII: Attack and defense speed (FPS). Settings: $m=2,\lambda=0.25,\Delta_{i}=\Delta_{o}=0.5$.
@@
Overhead Analysis. We analyze the computational efficiency on a single NVIDIA RTX A5000 GPU using the V2X-Sim test set under the default attack setting ($m=2,\lambda=0.25,\Delta_{i}=\Delta_{o}=0.5$). As detailed in Table VI, the average total defense latency per frame is around 26.0 ms, which is dominated by the temporal LSTM-AE inference. This latency corresponds to a processing speed of around 38.6 FPS, as compared with other methods in Table VII. Since this speed significantly exceeds the typical 10-20 Hz sensor frequency (e.g., LiDAR), GCP can verify collaborative messages within the generation cycle. This ensures a 0-frame detection delay (where attacks at frame $t$ are identified immediately), preventing perception bottlenecks. Furthermore, GCP is resource-efficient, incurring zero bandwidth overhead and negligible GPU memory cost (+118 MB), making it ideal for resource-constrained deployment. For extreme constraints, further optimizations like model distillation or FP16 quantization could be applied.
Generalizability to Different CP Model. We extend the evaluation to include DiscoNet [19] and When2com [21] besides the default V2VNet. DiscoNet utilizes distilled collaboration while When2com employing handshake communication for collaboration. Table VIII (averaged over three attack modes) shows that GCP consistently protects these diverse CP models, demonstrating its broad applicability and robustness across different collaborative fusion frameworks.
V-C Ablation Studies
Ablation on Defense Components. We evaluate the contributions of our defense components. First, we compare GCP with its spatial-only (GCP-S) and temporal-only (GCP-T) variants. As shown in Table IV, both components contribute to the overall defense. For PGD attacks, the spatial component is dominant, with GCP-S achieving 77.21% AP@0.5, comparable to the full model. However, for the challenging BAC attack, neither component alone suffices, with significant drops in GCP-S (7.41%) and GCP-T (10.63%). This confirms the necessity of our joint spatial-temporal design. Notably, GCP-S alone outperforms ROBOSAC [17] and MADE [35], validating our CSCLoss design.
Ablation on Hyper-Parameters. We analyze the frame cache length ($K$) and consecutive KF interpolation times ($L$). As shown in Fig. 6, moderate values yield optimal performance. The cache size peaks around $K=5$ (AP@0.5 of 76.87% under BAC S-mode), balancing context and noise. For interpolation times, performance is optimal at $L=3$ (77.42% AP@0.5 for BIM), as excessive interpolation may introduce cumulative errors.
Figure 9: Statistical dependency analysis and FDR control validation. It validates the PRDS assumption via positive correlations and demonstrates the effectiveness of the chosen operating point.
TABLE VIII: Generalizability of GCP across different CP models.
CP Model Method AP@0.5 AP@0.7
No Def. GCP No Def. GCP
DiscoNet PGD 37.05 78.35 36.25 77.28
C&W 17.88 77.62 15.28 76.35
BIM 38.69 78.21 37.62 77.25
BAC 55.42 77.45 45.58 76.42
When2com PGD 23.65 49.65 22.05 46.85
C&W 11.45 49.18 9.32 46.28
BIM 24.68 49.52 22.88 46.82
BAC 35.32 49.08 27.65 46.35
V-D Qualitative Results